# Multi-step Problem Solving Through a Verifier: An Empirical Analysis on Model-induced Process Supervision

### v1

@@ 4 Analysis
Figure 4: We apply the trained output- and process-supervised verifiers on various combinations of model and datasets. No verification and self consistency are also shown, wherever applicable. We use the default training objective that directly learns the estimate accuracies and the max aggregation function for the process verifier. In the x-axis, we vary the number of generated solutions to apply the verifier on, and in the y-axis we plot the performance (accuracy %). The standard deviation is also given. As a reference, we note that the purple line, representing the average performance of the generated solutions of the reasoner without any verification, matches the expectation to be an (almost) flat horizontal line with decreasing standard deviation. *While the reasoner that generates MiPS data and the reasoner that the verifier validates on is PaLM 2-L, the verifier is trained from a PaLM 2-S.
@@ 4.1 Models
In our experiments, we consider two LLMs, PaLM 2-S and PaLM 2-L Anil et al. (2023) to conduct our experiments on. We intend to understand the capability of MiPS data and analyze design choices of the verifier when trained on it. A concurrent work Wang et al. (2023) conducted a similar experiment on a different set of LLMs, namely LLama2, LLemma, Mixtral, and Deepseek Touvron et al. (2023); Azerbayev et al. (2023); Jiang et al. (2024); Bi et al. (2024).
@@ 4.2 Datasets
Table 1 contains detailed statistics about the datasets.
4.3 Settings
Throughout the paper, we choose to use a temperature value of $r_{g}=r_{mc}=0.7$ for both constructing MiPS and generating solutions on the test set. The number of generations for constructing MiPS $n_{g}$ is set to 32 for GSM8K and MBPP, and 8 for MATH. The number of completions $n_{mc}$ is also 32 for GSM8K and MBPP, and 8 for MATH. For GSM8K, we experiment with both PaLM 2-S and PaLM 2-L in the reasoner-verifier framework. For MATH, due to compute constraints, the reasoner we use is the PaLM 2-L, and the verifier trained is the PaLM 2-S. For MBPP, we find marginal differences in performance between using PaLM 2-S and PaLM 2-L as the reasoner, therefore we experiment only with the PaLM 2-S. During generation, all models are 8-bit quantized, and during training, we use a bfloat16 percision. Since MiPS contains an annotation for each intermediate step in the solution, it is naturally the number of steps times larger than output supervision. Therefore, we additionally generate more data to train the OSV. For training, we follow standard reward model training recipes, with an exception on the training epochs. Similar to Lightman et al. (2023), we also find it better to train the OSV for 1 epoch and the PSV for 2 epochs. For the OSV on MATH data, we find that training with a small 0.2 epochs (essentially training on less data) is better than training longer. For all experiments, we report the results of the average of 5 independently trained verifiers with different random seeds.
@@ 4.4 Directly Applying MiPS
Figure 6: We plot the performance of using various aggreagtion functions with PSV while restricting it to predict only the last $k$ steps or the last $p$ percentage of steps.
We first present the performance of the process verifier trained on MiPS data, using the default training objective on correctness scores directly, and the max aggregation function on the three datasets. For this experiment, we varied the number of solutions to be verified by the verifier from 2 to 128, to clearly depict the trend of the compared verifier’s performance. The results are shown in Fig. 4. The plots conveys several pieces of information.
@@
  • In all experiments, the verifier trained on MiPS using the max aggregation function showed stronger results than output verification. On GSM8K, the process verification is better than self consistency. On MATH, the performance lacks a bit. We note that this may be because we are training a less competent verifier (PaLM 2-S) than the reasoner (PaLM 2-L). Wang et al. (2023) showed improvements upon self consistency when the verifier and reasoner are of the same sizes.
  • The high performance of max may be unexpected if one believes max would be biased towards the first few correct steps of in in correct solution. In later analysis, we will show that (1) max favors high scores, similar to some other aggregation functions that perform well, that is preferred on MiPS data; (2) Due to higher noise in the earlier steps in MiPS data, the prediction scores of the earlier steps is of lower value (i.e., model confidence is lower), thus showing less affect to max.
  • In all experiments, the sum_logprob (product of probabilities) and min aggregation function are much worse than max and even using OSV, nevertheless still providing benefits over not using a verifier.
@@
From these results, we focus our analysis on two subjects: (1) the choice of aggregation functions, and (2) the affect of noise on generalization of the PSV.
@@ 4.5 Aggregation Functions
Table 2: We train a verifier based on PaLM 2-S and data generated by PaLM 2-S and test its applicability to transfer to validate solutions generated by two different reasoners, PaLM 2-L and gpt-turbo-3.5. A reference score of validing solutions generated by PaLM 2-S is also given. We evaluate this on both GSM8K and MBPP.
@@
To start with the analysis, we consider ten aggregation functions (sum and means of log probabilities, probabilities, logits, and odds, and maximum and minimum value over all steps). We obtain their performance on the MiPS dataset and plot it with the performance of the verifier using the aggragation function on the test set (Fig. 5). We first observe that in general, the two performances have positive correlations, indicating that it is possible to select a aggregation function on the MiPS dataset and use it during inference. Second, we notice that both min and sum_logprob have low performances not only during inference, but also in MiPS. This indicates that the poor performance of them likely is related to the construct of MiPS. Indeed, we realize that sum_logprob does not have a high correlation with a correct solution as it naturally penalizes long solutions. Similarly, for min, the possibility for a set of solutions to be wrongly verified using min is when the solution with the largest minimum correctness over all steps turns out to be wrong. This is not an unlikely event to happen, particularly in hard datasets where the largest minimum correctness is not high. To clarify these better, we answer the following questions:
@@
What are common in good aggregation functions for MiPS data? We believe a rule of thumb of a good aggregation function is a function that values highly scores highly. Consider two functions, one that values high scores (selects the solution with the highest high scores, e.g., max) and one that values the low scores (selects the solutions with the highest low scores, e.g., min). The first function is wrong only when the highest score solution is incorrect, in a simple case where there is only one observed step score for each solution, the probability is $1-s_{\text{max}},$ where $s_{\text{max}}$ is the score. Similarly, for min, the probability that the solution with the highest minimum score is wrong is $1-s_{\text{min}}.$ Since $s_{\text{max}}\geq s_{\text{min}},$ the first function shall be preferred. This is in line with the observation from Fig. 5 that aggregations of odds and logits are usually better than that of probabilities and log probabilities.
@@
Why did sum_logprob and min work well in Lightman et al. (2023) and Wang et al. (2023)? In Lightman et al. (2023), the dataset is constructed by human identifying all (earliest) incorrect steps, which corresponds to a prediction of 0 for the verifier (i.e., following the analogy in the previous discussion, this indicates that $s_{\text{max}}=s_{\text{min}}=1.0$). The min function would be correct on every instance in the training dataset, and if the verifier generalizes well, resembles human identification of mistakes on the test dataset. For Wang et al. (2023), we note two differences during MiPS data construction. First, their training objective is to predict the binary value of the correctness score. Second, they are using a different reasoner to complete the intermediate solutions than the reasoner that generates the solutions. We leave this difference for further investigations.
@@ 4.6 Different Length Aggregations
To understand how the intermediate steps in process supervision help the verifier model, we illustrate the result of applying an aggregation function to only the last $k$ steps or last $p$ percentage steps of the solutions in Fig. 6. In the upper three plots, we show the performance of an aggregation function sum_logit on the three datasets with $1\leq k\leq 5$. In the lower three plots, we show the performance of three aggregation functions on MATH with $10\leq p\leq 100$. We note that only the MATH dataset have solutions long enough such that looking at a percentage number of steps is sensible.
  • For all experiments, the performance increases with more steps considered from the end, up to some point. This indicates that the verifier predictions on the final few steps brings in increasing value, suggesting that process scores indeed is beneficial.
  • For most experiments, the performance starts to drop after including some earlier steps. This suggests that the quality of the predictions for the first steps are poor. We believe this is because MiPS has a poorer estimation of the first steps than the last steps, since (1) intuitively it is hard to predict the correctness of a very early solution even for humans, causing burden for the verifier to learn, and (2) Monte Carlo samplings that create MiPS data also tend to suffer from longer horizons.
@@
  • In all experiments, using the last-step process verifier predicted value is more beneficial than output supervision alone. Recall that this is not because of the problem of data quantity as we upscaled the data to train the output verifier. We suggest that this is because the process supervision data is of more diverse context, thus helping the model in generalization.
@@ 4.7 Transferring to a Different Reasoner
Finally, we provide an auxiliary experiment to check whether the trained verifiers would transfer to different reasoning models. We apply the verifiers trained on reasoning data generated by a PaLM 2-S and use it to valid solutions generated by stronger reasoners (reasoners having higher No Verifier accuracy). We find the sum_logit aggregation function working well in this case. The result is shown in Tab. 2, which shows that the trained verifier transfers to different and stronger reasoners with a strong validation ability, indicating that the verifier is not learning something overly specific to the reasoner that generates the data
@@ 5 Conclusion
In this work, we introduce MiPS to automatically annotate intermediate solutions for multi-step problem solving. Such data can be used to train a process supervised verifier that validates solutions generated by a reasoner. On two math datasets and one coding dataset, we demonstrated that MiPS improves the ability of picking the correct solution over an otherwise trained output supervised verifier. We conduct analysis on the aggregation function used to pick the solution and suggest that compared to verifiers trained on human-annotated process supervision, MiPS data trained verifiers prefer different aggregation functions. We also showed that such verifiers do not overly emphasize on the mistakes of the reasoner that produced the data, and can be transferred to different reasoners. Furture work could explore creating a scalable way to obtain MiPS data for each token in solutions to train a more competent verifier and use it to tune the reasoner via reinforcement learning.

### v2

@@ 4 Analysis
Figure 3: We apply the trained output- and process-supervised verifiers on various combinations of model and datasets. Self-consistency scores are given as a reference, however, it would be not applicable to general multi-step reasoning tasks (e.g., figure (d), coding). We use the default training objective that directly learns the estimated accuracies and the max aggregation function for the process verifier. In the x-axis, we vary the number of generated solutions to apply the verifier on, and in the y-axis we plot the performance (accuracy %). The standard deviation is also given. As a reference, we note that the purple line, representing the average performance of the generated solutions of the reasoner without any verification, matches the expectation to be an (almost) flat horizontal line with decreasing standard deviation. *While the reasoner that generates MiPS data and the reasoner that the verifier validates on is PaLM 2-L, the verifier is trained from a PaLM 2-S.
@@ 4.1 Models
In our experiments, we consider two LLMs, PaLM 2-S and PaLM 2-L Anil et al. (2023) to conduct our experiments on. We intend to understand the capability of MiPS data and analyze design choices of the verifier when trained on it. A concurrent work Wang et al. (2023) conducted a similar experiment on a different set of LLMs, namely LLama2, LLemma, Mixtral, and Deepseek Touvron et al. (2023); Azerbayev et al. (2023); Jiang et al. (2024); Bi et al. (2024). Detailed experimental settings and hyperparameters can be found in Appendix A.1.
@@ 4.2 Datasets
  • GSM8K Cobbe et al. (2021) is a dataset of grade school math problems.
  • MATH Hendrycks et al. (2021) is also a math word problems dataset. It consists of math problems of high school math competitions.
  • MBPP is an entry-level Python programming dataset. The questions are coding challenges along with a test case that defines the function format and the solutions are Python code that is expected to solve several hidden test cases.
Table 1 contains detailed statistics about the datasets, and Appendix A contains more information on how we split these datasets into training and evaluation.
@@ 4.3 Directly Applying MiPS
Figure 5: We plot the performance of using various aggregation functions with PSV while restricting it to predict only the last $k$ steps or the last $p$ percentage of steps.
We first present the performance of the process verifier trained on MiPS data, using the default training objective on correctness scores directly, and the max aggregation function on the three datasets. For this experiment, we varied the number of solutions to be verified by the verifier from 2 to 128, to clearly depict the trend of the compared verifier’s performance. The results are shown in Fig. 3. The plots convey several pieces of information.
@@
  • In all experiments, the verifier trained on MiPS using the max aggregation function showed stronger results than output verification. On GSM8K, the process verification is better than self-consistency. On MATH, the performance lacks a bit. We note that this may be because we are training a less competent verifier (PaLM 2-S) than the reasoner (PaLM 2-L). Wang et al. (2023) showed improvements upon self-consistency when the verifier and reasoner are of the same sizes.
  • The high performance of max may be unexpected, as max seemly would be biased towards the first few correct steps of an incorrect solution. In later analysis, we will show that (1) max favors high scores, similar to some other aggregation functions that perform well, that is preferred on MiPS data; (2) Due to higher noise in the earlier steps in MiPS data, the prediction scores of the earlier steps is of lower value (i.e., model confidence is lower), thus showing less affect to max.
  • In all experiments, the sum_logprob (product of probabilities) and min aggregation function are much worse than max or even using OSV, nevertheless still providing benefits over not using a verifier.
@@
From these results, we focus our analysis on two subjects: (1) the choice of aggregation functions, and (2) the effect of noise on generalization of the PSV.
@@ 4.4 Aggregation Functions
Table 2: We test on GSM8K the effect of different training objectives and aggregation functions (Hard &\&& min, the combination used in Wang et al. (2023), and Soft &\&& max, the combination we suggest). All 3 base models are 7B in size, and MM denotes the MetaMath Yu et al. (2023b) fine-tuned version of them.
Model Llemma MM-Llemma MM-Mistral
Soft + max 54.7 72.4 80.3
Soft + min 51.2 70.1 77.8
Hard + max 50.2 68.9 78.1
Hard + min 52.4 70.8 79.2
Table 3: We train a verifier based on PaLM 2-S and data generated by PaLM 2-S and test its applicability to transfer to validate solutions generated by two different reasoners, PaLM 2-L and gpt-turbo-3.5. A reference score of validating solutions generated by PaLM 2-S itself is also given. We evaluate this on two tasks, GSM8K and MBPP.
@@
To start with the analysis, we consider ten aggregation functions (sum and means of log probabilities, probabilities, logits, and odds, and maximum and minimum value over all steps). We obtain their performance on the MiPS dataset and plot it with the performance of the verifier using the aggregation function on the test set (Fig. 4). We first observe that in general, the two performances have positive correlations, indicating that it is possible to select an aggregation function on the MiPS dataset and use it during inference. Second, we notice that both min and sum_logprob have low performances not only during inference, but also in MiPS. This indicates that the poor performance of them likely is related to the construct of MiPS. Indeed, we realize that sum_logprob does not have a high correlation with a correct solution, as it naturally penalizes long solutions. For min, the possibility for a set of solutions to be wrongly verified using min is when the solution with the largest minimum correctness over all steps turns out to be wrong. This is actually not an unlikely event to happen, particularly in consideration when the reasoner makes an erroneous continuation to an initially correct solution. To clarify these better, we answer the following questions:
@@
What are common in good aggregation functions for MiPS data? We believe a rule of thumb of a good aggregation function is a function that values high scores highly. Consider two functions, one that values high scores (selects the solution with the highest high scores, e.g., max) and one that values the low scores (selects the solutions with the highest low scores, e.g., min). The first function is wrong only when the highest score solution is incorrect, in a simple case where there is only one observed step score for each solution, the probability is $1-s_{\text{max}},$ where $s_{\text{max}}$ is the score. Similarly, for min, the probability that the solution with the highest minimum score is wrong is $1-s_{\text{min}}.$ Since $s_{\text{max}}\geq s_{\text{min}},$ the first function shall be preferred. This is in line with the observation from Fig. 4 that aggregations of odds and logits are usually better than that of probabilities and log probabilities.
@@
Why did sum_logprob and min work well in Lightman et al. (2023) and Wang et al. (2023)? In Lightman et al. (2023), the dataset is constructed by human identifying all (earliest) incorrect steps, which corresponds to a prediction of 0 for the verifier (i.e., following the analogy in the previous discussion, this indicates that $s_{\text{max}}=s_{\text{min}}=1.0$). The min function would be correct on every instance in the training dataset, and if the verifier generalizes well, resembles human identification of mistakes on the test dataset. For Wang et al. (2023), we note a difference during MiPS data construction as their training objective is to predict the binary value of the correctness score, we discuss this more in Sec 4.6.
@@ 4.5 Different Length Aggregations
To understand the generalization issue, we illustrate the result of applying an aggregation function to only the last $k$ steps or last $p$ percentage steps of the solutions in Fig. 5. In the upper three plots, we show the performance of an aggregation function sum_logit on the three datasets with $1\leq k\leq 5$. In the lower three plots, we show the performance of three aggregation functions on MATH with $10\leq p\leq 100$. We only conducted this analysis on the MATH dataset, as it have solutions long enough such that looking at a percentage number of steps is sensible.
  • For all experiments, the performance increases with a few more steps considered from the end. This indicates that the PSV predictions on the last steps brings in increasing value, suggesting that process scores indeed are beneficial.
  • For most experiments, the performance starts to drop after including some early steps. This suggests that the quality of the predictions for the first steps are poor. We believe this is because MiPS has a poorer estimation of the first steps than the last steps, since intuitively it is hard to predict the correctness of a very early solution, causing burden for the verifier to learn.
@@
  • In all experiments, using the last-step process verifier predicted value is more beneficial than output supervision alone. Recall that this is not because of the problem of data quantity, as we upscaled the data to train the output verifier. We suggest that this is because the process supervision data is of more diverse context, thus helping the model in generalization.
4.6 Different Training Objectives
The main difference in the method of ours and Wang et al. (2023) is the training objective of the verifier, where we train the verifier to directly predict the estimated accuracies (Soft Objective), and they train the verifier to predict a binarized value (non-zeroness) of the accuracy (Hard Objective). In our previous analysis, we noted that since the reasoner is imperfect, MiPS would provide underestimated accuracies of the intermediate steps, which is harmful to aggregation functions that focus on low values (e.g., min). In contrast, the non-zeroness of the accuracy would cause an overestimation of the accuracy, which, by the same argument, would be harmful to aggregation functions that focus on high values (e.g., max). To verify this, we conduct the experiments using the same language model as Wang et al. (2023) on the GSM8K dataset, using both training objectives and aggregation functions.
The experiment setting is detailed in Appendix A.2. The results are in Table 2. It is observed that, indeed, the max aggregation is better for the soft objective and the min aggregation is better for the hard objective. It also turns out that soft objective with the max aggregation consistently outperforms hard objective with min aggregation. We believe this to be a strong motivation for the use of the soft objective in MiPS.
@@ 4.7 Transferring to a Different Reasoner
Finally, we provide an auxiliary experiment to check whether the trained verifiers would transfer to different reasoning models. We apply the verifiers trained on reasoning data generated by a PaLM 2-S and use it to valid solutions generated by stronger reasoners (reasoners having higher No Verifier accuracy). We find the sum_logit aggregation function working well in this case. The result is shown in Tab. 3, which shows that the trained verifier transfers to different and stronger reasoners with a strong validation ability, indicating that the verifier is not learning something overly specific to the reasoner that generates the data.
@@ 5 Conclusion
In this work, we introduce MiPS to automatically annotate intermediate solutions for multi-step problem solving. Such data can be used to train a process supervised verifier that validates solutions generated by a reasoner. On two math datasets and one coding dataset, we demonstrated that MiPS improves the ability of picking the correct solution over an otherwise trained output supervised verifier. We conduct analysis on the aggregation function used to pick the solution and suggest that compared to verifiers trained on human-annotated process supervision, MiPS data trained verifiers prefer different aggregation functions. We also showed that such verifiers do not overly emphasize on the mistakes of the reasoner that produced the data, and can be transferred to different reasoners. Future work could explore creating a scalable way to obtain MiPS data for each token in solutions to train a more competent verifier and use it to tune the reasoner via reinforcement learning.
6 Limitation
6.1 Underperformance on the MATH dataset
In our work, we did not manage to conduct all experiments using the same, large model. Especially for the MATH dataset, we had to train a smaller verifier to compensate of the long sequence length and data size. This probably led to us finding a lower performance of process and output verifier than the straightforward self-consistency. We believe in general that this is not true, as Lightman et al. (2023) and Wang et al. (2023) both showed that process/output verifier should output self-consistency on the MATH dataset.
6.2 Efficiency
MiPS, while automatic, requires a non-trivial amount of computation effort in generating the dataset to train the verifier. We did not attempt to reduce the computational effort, as we’d like to show the most direct comparison with no verifiers and output supervised verifiers. We do believe it is very possible to reduce the computation costs, for example, by avoiding creating data on every intermediate solutions, and we suggest future work to explore this direction.
Appendix A Datasets
A.1 Settings
Throughout the paper, we choose to use a temperature value of $r_{g}=r_{mc}=0.7$ for both constructing MiPS and generating solutions on the test set. The number of generations for constructing MiPS $n_{g}$ is set to 32 for GSM8K and MBPP, and 8 for MATH. The number of completions $n_{mc}$ is also 32 for GSM8K and MBPP, and 8 for MATH. For GSM8K, we experiment with both PaLM 2-S and PaLM 2-L in the reasoner-verifier framework. For MATH, due to compute constraints, the reasoner we use is the PaLM 2-L, and the verifier trained is the PaLM 2-S. For MBPP, we find marginal differences in performance between using PaLM 2-S and PaLM 2-L as the reasoner, therefore we experiment only with the PaLM 2-S. During generation, all models are 8-bit quantized, and during training, we use a bfloat16 precision. Since MiPS contains an annotation for each intermediate step in the solution, it is naturally the number of steps times larger than output supervision. Therefore, we additionally generate more data to train the OSV. For training, we follow standard reward model training recipes, with an exception on the training epochs. Similar to Lightman et al. (2023), we also find it better to train the OSV for 1 epoch and the PSV for 2 epochs. For the OSV on MATH data, we find that training with a small 0.2 epochs (essentially training on less data) is better than training longer. For all experiments, we report the results of the average of 5 independently trained verifiers with different random seeds.
A.2 Settings of the Objective Experiment
To reduce the cost, when conducting the experiment to compare the two training objectives, we scaled down the experiment. On GSM8K, we used 2000 data points for verification training data generation, and $n_{g}=n_{mc}=8$.